Steven is planning a big garden that is 25 meters longer than its width. The garden will have an area of 5100m^2.
Find the dimensions of the garden.

Answer by mananth(16946)

let the width be x
the length will be x+25
Area = l*w
x(x+25) = 5100
x^2+25x-5100=0
x^2+85x-65x-5100=0
x(x+85)-60(x+85)=0
(x-60)(x+85)=0
x= 60
width = 60
length = 60+25= 85 m